Report of extended tehnical stop activity and barrel hw alignment
This presentation is the property of its rightful owner.
Sponsored Links
1 / 31

Report of extended tehnical stop activity and Barrel HW alignment PowerPoint PPT Presentation


  • 86 Views
  • Uploaded on
  • Presentation posted in: General

Report of extended tehnical stop activity and Barrel HW alignment. Szillási Zoltán on behalf of the HW Muon Alignment Group. (ATOMKI,Debrecen). Muon Barrel Workshop, February 24, 2011. This project is supported by the Hungarian National Research Fund (OTKA) No.: NK67974. Outline.

Download Presentation

Report of extended tehnical stop activity and Barrel HW alignment

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Report of extended tehnical stop activity and barrel hw alignment

Report of extended tehnical stop activity andBarrel HW alignment

Szillási Zoltán on behalf of the HW Muon Alignment Group

(ATOMKI,Debrecen)

Muon Barrel Workshop, February 24, 2011.

This project is supportedbytheHungarian National Research Fund (OTKA) No.: NK67974


Outline

Outline

  • LinkDisk repositioning

  • Z-study of the Barrel

  • (very) preliminary data on the startup barrel geometry based on HW Ali

2011.02.24.

Muon Barrel WorkShop2


Report of extended tehnical stop activity and barrel hw alignment

LinkDiskrealignment


Report of extended tehnical stop activity and barrel hw alignment

HW Interventions in January 2011 (from Gervasio)

  • Jan 6:

    • LV problems in both endcaps until 15.30

      • LOTO problem: cause/person still unknown… under investigation from LV team

    • Adjust LD-

      • Four AR lines get to YE-1 TPs

    • Fixed two link electronic boards

      • One ELMB for BD- and one Power Card in YB-2 for line -315

  • Jan 7:

    • Test insertion of ½ T1: near-side

    • 10:30 LD- moves again!

      • See next slide

      • LD- has unwanted friction

    • Test SLM lines in YE-: everything OK

    • 16:30 Realign LD- again

    • Loose LV: no more readout

2011.02.24.

Muon Barrel WorkShop4


Report of extended tehnical stop activity and barrel hw alignment

HW Interventions in January 2011

  • Jan 7:

    • REALIGNED LD- (a second time) in the afternoon

  • Jan 8:

    • Test SLM lines again in YE-: everything OK

  • Jan 10:

    • No LV, CMS wide: no more readout

  • Jan 14-18:

    • T1+ installation

    • NO LV, no readout of LD+ or SLM+

2011.02.24.

Muon Barrel WorkShop5


Report of extended tehnical stop activity and barrel hw alignment

State after Jan 7

  • LD-

    • Four AR lasers reach YE-1 after final adjustment Jan 7

  • Link electronics

    • Apparently both fixes went OK: no more crashes so far

  • SLM-

    • All SLMs go across both both T1 halves

      • Line SLM 1-4 in YE-3 was temporarily blocked, so this one not really tested but suspect OK

  • LD+ and SLM+

    • All T1 installation happened with no LV: no readout

      • CMS chose to take this risk a few months back

    • Can only hope we will find everything OK when LV is restored

Positive side has been tested since and it is OK

2011.02.24.

Muon Barrel WorkShop6


Report of extended tehnical stop activity and barrel hw alignment

State NOW

  • LD-

    • Four AR- lasers reach YE-1 at 0 T

    • Four AR- lasers reach YE-1 at 3.8 T !!!

    • Reco of Link at both ends now possible at 0T and 3.8T

      • See status talk

  • Link electronics

    • ELMB in YB0 still causes problems (we need to run continuously for several days for the card to send errors)

      • This is completely transparent for Link: our data taking script simply reinitializes the card when this happens with zero impact on link data

    • The YB-2 problem was serious but this was fixed fine

2011.02.24.

Muon Barrel WorkShop7


Report of extended tehnical stop activity and barrel hw alignment

AR- lasers  YE-1 @ 3.8T

2011.02.24.

Muon Barrel WorkShop8


Report of extended tehnical stop activity and barrel hw alignment

LD- readjustment

(from Celso)

  • Since yesterday (2011.02.10) we are at Full Field back again: B=3.8T

    • Our distancemeters indicate a good closing of the detector:

Asymmetry after incident

Symmetry recovered!

2011.02.24.

Muon Barrel WorkShop9


Report of extended tehnical stop activity and barrel hw alignment

Z-study of the Barrel YB±2 wheelsafterthe magnet ramp

(Noemi Beni)


Z movement of yb 2 2011 long

Z movement of YB±2 2011 long

Movements of the Z-bar LEDs in the Z direction.

YB-2

YB+2

1.5mm

2mm

15°

15°

Ref 3888 @0T (range 2011. Feb 03-21)

1.1mm

1.5mm

75°

75°

1.5mm

1.5mm

135°

135°

1.5mm

2mm

195°

195°

2.6mm

2.4mm

255°

255°

0T

0T

0T

0T

2.5mm

2.1mm

1T

1T

2T

2T

315°

315°

3.8T

3.8T

Something new!!!

2011.02.24.

Muon Barrel WorkShop11


Phi movement of yb 2 2011 long

Phi movement of YB±2 2011 long

Movements of the Z-bar LEDs in the Phi direction.

YB-2

YB+2

0.1mm

0.2mm

15°

15°

0mm

Ref 3888 @0T (range 2011. Feb 03-21)

0.2mm

75°

75°

0.2mm

0.1mm

135°

135°

0.3mm

0.1mm

195°

195°

0.1mm

0.5mm

255°

255°

0T

0T

0T

0T

0.4mm

0.5mm

1T

1T

315°

315°

2T

3.8T

2T

3.8T

2011.02.24.

Muon Barrel WorkShop12


Report of extended tehnical stop activity and barrel hw alignment

Z-movement of YB±2 are compatible with the last magnet cycle

no long term (~2months) relaxation effect is seen

2011.02.24.

Muon Barrel WorkShop13


Report of extended tehnical stop activity and barrel hw alignment

Phi-movement of YB±2 are compatible with the last magnet cycle

no long term (~2months) relaxation effect is seen

2011.02.24.

Muon Barrel WorkShop14


Phi movement of yb 2 2010 2011

Phi movement of YB±2 2010 & 2011

2011 first & 2010 last

2010 first

Interpretation of the measured data:

75°

75°

135°

135°

15°

15°

YB+2

YB-2

195°

195°

315°

315°

255°

255°

  • from Phi displacement plot

  • the arrows don’t show absolute values

  • the arrows are the same for the last magnet ramp up in 2010 and 2011

2011.02.24.

Muon Barrel WorkShop15


Z movement 2010 2011

Z movement 2010 & 2011

2011 first & 2010 last

2010 first

Interpretation of the measured data:

ViewfromNearside

75°

YE+

YE-

YE-

YE+

YB+2

YB+1

YB0

YB-1

YB-2

15°

315°

Viewfrom Far side

YB+2

YB+1

YB0

YB-1

YB-2

135°

195°

  • from Z displacement plot

  • the arrows don’t show absolute values

  • the arrows are the same for the last magnet ramp up in 2010 and 2011

255°

2011.02.24.

Muon Barrel WorkShop16


Report of extended tehnical stop activity and barrel hw alignment

(very) preliminarydataontheinitial 3.8T geometry of the Barrel


Missing chambers

”Missing” chambers

New dataset from run No 3981 (2011.02.13. 10:36-13:02) has been choosen.

Measurement is coordinated with the Link and EndCap

Only qualitative results yet!

2011.02.24.

Muon Barrel WorkShop18


3981 vs 2800 blessed

3981 vs 2800 (blessed)

(thanks to Aysen!!!)

Translation magn 200x

Rotation magn 200x

New problem! To be checked

Something

in 2800?

Due to a missing MAB

Known probem (can be cured)!

2011.02.24.

Muon Barrel WorkShop19


3981 vs 2800 blessed1

3981 vs 2800 (blessed)

YB-2

YB-1

YB0

Translation magn x200

Rotation magn x200

YB+2

YB+1

Missing MAB +1/9


Plans

Plans

  • Finalize 3981 asap and bless

  • Comparision of relative movements of 3.8T-0T wrt Survey is ongoing – our results are foreseen in a week

  • APE analysis started (see Gervasio’s talk)

2011.02.24.

Muon Barrel WorkShop21


Conclusions

Conclusions

  • LinkDisk repositioning went well

  • Z-study of the Barrel – no long term relaxation seen

  • (very) preliminary data on the startup barrel geometry based on HW Ali – minor differences seen wrt. 2800


Report of extended tehnical stop activity and barrel hw alignment

Backup slides


Report of extended tehnical stop activity and barrel hw alignment

Recall of blessing 2800


Missing chambers1

”Missing” chambers

Wheel/Sector/Station

  • -2/11/4

  • -1/13/4

  • -0/14/4

  • +1/7/4 Can’t measure any LED (communication problem with chamber).

  • +2/1/2 Can’t measure any LED (communication problem with chamber).

  • +2/1/4

  • +2/2/4

  • +2/4/4

    The following are left uncorrected due to bad fit from Luca:

  • Wheel -2 Sector 5 MB4

  • Wheel 0 Sector 13 MB4


Comparison 2800 vs 1125

Comparison 2800 vs 1125

(thanks to Aysen!!!)

Translation magn 200x

Rotation magn 200x

Real thing?

Due to a missing MAB

Known probem!

2011.02.24.

Muon Barrel WorkShop26


2800 vs 1125

2800 vs 1125

YB-2

YB-1

YB0

Translation magn x200

Rotation magn x200

Measured only from one side

YB+2

YB+1

Missing MAB +1/9


Strange chambers

Strange chambers

In RPhi or Z

  • Wheel +2 Sector 11 MB4

    • Problem with the camera calibration: +2/10 4EXT L, CamId: 81, SensorAngY=3.78 deg

  • Wheel 0 Sector 10 MB1Fork 3 6side Fork 4 4side are missing

  • Wheel 0 Sector 10 MB3Fork 3 6side Fork 4 4side are missing

  • Wheel +1 Sector 10 MB2 Fork 1 6side, Fork 2 6side 4LED, F3 6side, F4 4 side are missing

  • Wheel +1 Sector 10 MB4 Fork 1, F2, F3 6side, F4 4side are missing

  • Wheel +1 Sector 1 MB3Fork3 6side and Fork4 4side are missing, Fork4 6side bad

  • Wheel -1 Sector 1 MB1 Fork 2 6 side 3 LED missing

  • Wheel -1 Sector 3 MB4 Fork 1 6side 2 LED, 4side 2LED are missing, F2 bad

  • Wheel -1 Sector 3 MB3 Fork 1 4side 2 LED missing, F2 bad

    In R or rootZ (global)

  • Wheel -1 Sector 1 MB3 (chimney chamber)Fork 1 6side, Fork 2 missing

  • Wheel 0 Sector 2 MB1 Fork 2 is bad, (see next slides)

2011.02.24.

Muon Barrel WorkShop28


Comparison 2800 vs 11251

Comparison 2800 vs 1125

Translation magn 1000x

Rotation magn 1000x

This is just to show that there are micro movements…

2011.02.24.

Muon Barrel WorkShop29


2800 vs 11251

2800 vs 1125

YB-2

YB-1

YB0

Translation magn x1000

Rotation magn x1000

YB+2

YB+1


Z bar measurements on yb 2

Z-bar measurements on YB±2

What camera observes:

Camera local X

‘Real’ life:

IP

ZLED1

ZLED4

Camera local Y

+

ZLED2

ZLED5

ZLED1

ZLED4

+

ZLED3

ZLED6

ZLED2

ZLED5

YB moves (with MAB) instead of Zfork(changes sign in φ, but not in Z due to an other trafo.)

ZLED6

ZLED3

+φCMS

on posZ side

+φCMS

on negZ side

2011.02.24.

Muon Barrel WorkShop31


  • Login